LIF SPECTROSCOPY OF A Σ − Σ TYPE BAND SYSTEM IN THE C_{2}H_{2}/ CS_{2} DISCHARGE PLASMA

\underline{MASAKAZU NAKAJIMA}^{a}, HITOMI TOYOSHIMA, YU YONEDA, YOSHIHIRO SUMIYOSHI, and YASUKI ENDO, Department of Basic Science, Graduate School of Arts and Sciences, The University of Tokyo, Tokyo 153-8902, Japan.

New vibronic band system has been observed in the 21800−23000 cm^{−1} region by laser-induced fluorescence spectroscopy in a discharged supersonic jet of the C_{2}H_{2} and CS_{2} mixture gas. Several chemical tests indicate that the spectral carrier contains H, C and S atom(s). High-resolution spectra of these vibronic bands show Σ − Σ type rotational structures. The origin band position and the effective rotational constants of the upper and lower levels have been determined to be T_{0}=21877.7706(7), B′_{eff} =0.04932(1) and B′′_{eff} =0.05013(2)cm^{−1}, respectively. Based on the results of ab initio calculations, it is considered that bent HSCCS is the most probable candidate for the spectral carrier.

aCurrent address: Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan